# Causal reasoning in difference graphs

### v1

@@ 5 Related works
There has been extensive research on causal reasoning using graphs that represent families of graphs. For example, Maathuis and Colombo [2013], Perkovic et al. [2016], Perkovic [2020], Jaber et al. [2022], Wang et al. [2023] extended causal reasoning to partially oriented graphs such as CPDAGs and PAGs. PAGs are similar to CPDAGs but allow for hidden confounding. Both CPDAGs and PAGs represent multiple causal DAGs that share the same skeleton, some common edge orientations, and the same topological ordering. From a different perspective, Iwasaki and Simon [1994], Chalupka et al. [2016], Anand et al. [2023], Assaad et al. [2023], Ferreira and Assaad [2024], Assaad et al. [2024] have examined causal reasoning using cluster graphs, where each vertex represents a cluster of low-level variables. These cluster graphs represent various causal DAGs, with each vertex corresponding to a different low-level variable. Most of these works assumed that the causal DAGs represented by the cluster graphs share the same topological ordering, though some relaxed this assumption. In addition, other works, such as Richardson [1997], Forré and Mooij [2020], have specifically focused on cyclic directed graphs, where vertices do not represent clusters. However, to date, there has been no investigation into causal reasoning using difference graphs. The existing work in the literature cannot be directly applied to difference graphs because the absence of an edge in these graphs has a different interpretation than in other types of graphs.
@@ 6 Conclusion
This paper is the first to explore causal reasoning through the lens of difference graphs. While difference graphs may not be as intuitive as classical causal DAGs for identifying causal effects, this work establishes conditions for identifying total effects in a nonparametric setting and direct effects in a linear setting, directly from the difference graph. These findings also demonstrate that difference graphs can be effectively applied to identify causal changes—specifically, total and direct causal changes—providing a valuable tool for causal inference and root cause analysis in complex systems. For future works, it would be valuable to explore the controlled direct effect and natural direct effect in a nonparametric setting. Additionally, providing conditions for the identifiability of path-specific effects using difference graphs would be an important advancement. This paper focuses on cases where the difference graph is fully specified. However, some algorithms detect only partially oriented difference graphs, making it worthwhile to extend this work to such partially specified graphs. Another open question is the utility of discovering differences in the presence of hidden confounding in real-world applications, as the precise definition of this remains unclear. If it is shown that difference graphs with hidden confounding are useful and can be learned or provided by experts, it will be crucial to investigate the identifiability conditions for these graphs as well.
@@ Acknowledgements
This work was supported by the CIPHOD project (ANR-23-CPJ1-0212-01).

### v2

@@ 5 Related works and discussion
There has been extensive research on causal reasoning using graphs that represent a class of graphs. For example, Maathuis and Colombo [2013], Perkovic et al. [2016], Perkovic [2020], Jaber et al. [2022], Wang et al. [2023] extended causal reasoning to partially oriented graphs such as CPDAGs and PAGs. PAGs are similar to CPDAGs but allow for hidden confounding. Both CPDAGs and PAGs represent multiple causal DAGs that share the same skeleton, some common edge orientations, and the same topological ordering. Moreover, extensive research has been conducted for recovering these structures from observational data under specific assumptions [Spirtes et al., 2000, Glymour et al., 2019]. From a different perspective, Iwasaki and Simon [1994], Chalupka et al. [2016], Anand et al. [2023], Assaad et al. [2023], Ferreira and Assaad [2024], Assaad et al. [2024] have examined causal reasoning using cluster graphs, where each vertex corresponds to a cluster of low-level variables, and each edge denotes a causal relationship between two clusters. A cluster graph can represent multiple causal DAGs but does not necessarily share the same skeleton as the true causal DAG and is not necessarily acyclic. Like CPDAGs and PAGs, cluster graphs can also be discovered from data under certain assumptions [Wahl et al., 2024]. In addition, other works, such as Richardson [1997], Forré and Mooij [2020], have specifically focused on cyclic directed graphs, where vertices do not represent clusters. More closely related to difference graphs are selection diagrams [Bareinboim and Pearl, 2012] which can also represent differences between populations. There is a substantial body of literature on their use in causal reasoning [Bareinboim and Pearl, 2012] and methods for discovering them from data [Li et al., 2023]. However, these two graphical representations serve different purposes. Selection diagrams provide more detailed information about stable causal relations and changes in noise, whereas difference graphs are more informative regarding changes in direct effects. In summary, to date, there has been no investigation into causal reasoning using difference graphs. The existing work in the literature cannot be directly applied to difference graphs because the absence of an edge in these graphs has a different interpretation than in other types of graphs.
This work addresses this gap and allow analysts to focus solely on the differences between two populations, simplifying reasoning, while still achieving identification for each domain individually. I emphasize that in this work, I assume that the difference graph is given and correct. The difference graph may be provided either by domain experts or by an algorithmic discovery method. In many domains, experts cannot construct a difference graph without knowing the true causal DAGs for both populations. However, in certain cases, this is feasible. For example, consider two cell populations: Population A (untreated) and Population B (treated with a drug that inhibits a specific protein). While the complete causal DAGs of these populations may be complex and unknown, domain knowledge about the drug’s direct mechanism of action allows researchers to construct a difference graph representing the expected changes in causal structure. When domain experts cannot directly specify a difference graph, they may rely on algorithmic approaches to discover it from data. Some algorithms discover an equivalence class of graphs [Wang et al., 2018, Bystrova et al., 2024], while others aim to recover the exact difference graph [Chen et al., 2023]. If the exact difference graph is known, all theorems intorduced in this paper, can be applied directly. For instance, consider the difference graph shown in [Chen et al., 2023, Figure 4 (a)], discovered from a real ovarian cancer dataset. Based on Theorem 1, the total effect of BIRC3 on PRKAR2B is not identifiable, whereas according to Theorem 3, the direct effect of BIRC3 on PRKAR2B is identifiable, assuming linearity. In contrast, when working with Markov equivalence classes of difference graphs, the applicability of the results of this paper depends on whether the conditions of theorems hold across all graphs in the equivalence class. For instance, consider a scenario with four variables: $X$, $Y$, $W_{1}$, and $W_{2}$. If every difference graph within a given Markov equivalence class includes the edges $X\leftarrow W_{1}$, $X\rightarrow W_{2}$, and $X\rightarrow Y$, then the conditions of Theorem 1 are satisfied across the entire equivalence class, ensuring that the total effect of $X$ on $Y$ is identifiable. Similarly, for the identification of the direct effect, suppose that in every difference graph within a given Markov equivalence class, the edges $W_{1}\leftarrow Y$, $W_{2}\rightarrow Y$, and $X\rightarrow Y$ are present. This guarantees that the conditions of Theorem 3 hold, allowing for the identification of $\alpha_{x,y}$.
@@ 6 Conclusion
This paper is the first to explore causal reasoning through the lens of difference graphs. While difference graphs may not be as intuitive as classical causal DAGs for identifying causal effects, this work establishes conditions for identifying total effects (and total causal changes) in a nonparametric setting and direct effects (and direct causal changes) in a linear setting, directly from the difference graph. For future works, it would be valuable to extend the results of this paper to direct effects in a nonparametric setting to path-specific effects. Additionaly, if it will be shown that difference graphs with hidden confounding are useful and can be learned or provided by experts, then it will be valuable to investigate the identifiability conditions in the presence of hidden confounding.
@@ Acknowledgements
I thank Elias Bareinboim from Columbia University for the insightful discussions on difference graphs and selection diagrams. This work was supported by the CIPHOD project (ANR-23-CPJ1-0212-01).